# From Voice to Safety: Language AI Powered Pilot-ATC Communication Understanding for Airport Surface Movement Collision Risk Assessment

### v1

@@ 3.2 Airport Surface Collision Risk Model
Mathematically, the collision risk is calculated based on the joint probability that each aircraft reaches the same area at the same timestamp of the node-link airport layout graph. The location of a potential collision can be the intersection of runways, taxiways, taxilanes, or the combination of any two. The collision risk model is defined as the joint distribution of the time overlap for two aircraft occupying the same area around a node in the airport layout map. In this section, we provide the the full probabilistic formulation of collision risk at a certain node in the node-link graph.
@@ 3.2.1 Total Travel Time Modeling
The -th aircraft travels a total of taxiway links until reaching the area of interest (i.e., potential collision spot), where the total travel time is given by . We assume each taxiway links has an associated distance and a taxi speed that is log-normally distributed with parameters and , which is,
@@
In practice, we approximate for any time by either Monte Carlo Simulations or Moment-Matching Approximations. For the convolution of log-normal distributions with moderate variance and , the Fenton-Wilkinson approach provides a feasible solution to directly match the first two moments, and is widely-adopted as the approximated analytical solution of log-normal sums in various field [115, 116, 117]. Fenton found that the sum of several independent log-normal variables can be reasonably approximated by another log-normal. Under this approximation, the mean of the route travel time equals the sum of link means, and the variance equals the sum of all link variance and covariance terms. These matched moments define the parameters of an approximate log-normal for the route [115]. Furthermore, [87] note that this log-normal approximation is computationally efficient compared to brute-force convolution or simulation, with only a modest loss of accuracy. That is, we are looking for the parameters of an approximate distribution of , where . That is,
and the associated cumulative density function (CDF) as,
@@
3.2.2 Risk Formulation
As reviewed in Section 2.2, the collision between two moving aircraft on the surface is quantified by the probability of conflicts between aircraft by evaluating random deviations in position and speed. This also guides our spatiotemporal risk formulation. In our formulation, we assume that a collision occurs when two interchangeable aircraft are satisfying the following two conditions,
@@ 3.2.2 Risk Formulation
We derive an expression to approximate the probability a collision occurs at at any time . We start with a joint probability distribution for the spatiotemporal state of the aircraft. Define be the conditional PDF of aircraft 1’s arrival time at location , be the PDF of aircraft 2’s arrival time location , and be the PDF describing aircraft 2’s spatial position at time .
@@
Approximating the velocity at collision as independent of and , a change of variables from from space to time near can be made using the expected inverse of the speed at the point of collision,
@@
We prove Equation 17 using the joint density of the position, arrival time, and velocity and the relation . We assume that the joint distribution of position and arrival time is independent of the aircraft velocity at the potential collision location as follows,
We first use Bayes’ rule for conditional distributions to relate the conditional position-velocity distribution to the conditional arrival time distribution, then utilize the independence of the velocity distribution.
The relation is utilized for a change of variables in the denominator spatial integral to cancel out the numerator temporal integral.
The position conditional distribution can be expressed in terms of a velocity integral of the temporal conditional distribution expression, resulting in the term.
@@ 4 Case Studies
In this section, we present two illustrative case studies to demonstrate our proposed framework. The first case study examines the 2024 Haneda Airport runway collision in Section 4.1, while the second focuses on the 2024 KATL taxiway collision accident in Section 4.2. Both accidents underscore the complex interplay of human error, communication misunderstandings, and operational failures. However, we use the assumed values for the link travel time speed distribution parameters of the first case study, based on recommended values mentioned in the FAA airport design recommendations, 150/5300-13B, as in [15]. For the second study, we obtain the ground movement data (i.e., ASDE-X) from Sherlock and use the link travel time speed distribution parameters from the real-world data. Details is given in Section 4.2.1.
4.1 Henada Airport Runway Incursion
The Haneda Airport collision occurred on January 2, 2024, when Japan Airlines Flight 516 (Airbus A350) collide with a Japan Coast Guard DHC-8-315Q aircraft on the runway. The Coast Guard aircraft was stationed on the runway to deliver relief supplies following the Noto Peninsula earthquake. Despite the dramatic collision and ensuing fire, all 379 occupants aboard the A350 were successfully evacuated, while five of the six crew members on the smaller aircraft lost their lives. Investigations have attributed the accident primarily to miscommunication and human error where the Coast Guard pilot misinterpreted air traffic control instructions and mistakenly believed he had clearance to enter the runway. The communication transcripts ahead of the occurrence of the accident was released and tested against our developed learning model as in Figure 7(a) and Figure 7(b). It is clear that JA722A was given clearance to hold at C5 and should wait for departure clearance. However, the JA722A goes to the runway and misunderstood as the first to take off at runway 34R. Our ATC Rule-Enhanced Learning modele processed the timeline of clearance given from the controller and the state of each aircraft and listed in Table 2.
@@ 4.1 Henada Airport Runway Incursion
Based on Table 2, we build the surface movement simulation environment with the airport node-link graph layout obtained from NASA FACET [23]. We simulate the occurrence of the accident from the timestamp where the last clearance was given to JA722A from the ATCo in Figure 8. The collision happened at the timestamp of the lower left figure. The simulation takes three inputs for each link, the link distance , the link travel speed parameters . was calculated based on the location of the node on the node-link graph, with the speed parameters assumed to be if the link is between runway nodes, and between runway and taxiways. For links between two taxiway nodes, we assume the speed follows , and for all other scenarios (taxilanes, ramps, etc). These speed assumption are based on the recommended values mentioned in the FAA airport design recommendations, 150/5300-13B, as in [15]. A detailed explanation of taxiways, runways, and terminologies are given in Figure 9.
Table 2 shows that the NER model is able to capture the destination runway for both Japan Air 516 and Delta 276, where the key information such as callsign, aircraft state and also retrieved. The destination node from the node-link graph are correlated by either, (a) the entry node of the given runway; (b) the similarity between the destination node name and the entities from pilot-ATC communication transcripts. Table 2 visualizes several timestamps of the simulation, based on the assumed speed distribution parameters. Incorporating the information retrieved from the NER model, this simulation successfully replicates the occurrence of the Henada runway incursion accident. The associate risk at each target node location is given in Figure 13(a), where node has the highest collision risk probability, matching the real-world crash.
4.2 KATL Taxiway Collision
In the second case study, we investigate the taxiway collision happened at KATL on September 10, 2024, where two Delta Air Lines aircraft collided while taxiing. An Airbus A350, preparing for an international departure to Tokyo, struck the tail of an Endeavor Air Bombardier CRJ-900, which was scheduled for a domestic flight to Lafayette, Louisiana. The collision, which occurred at the intersection of two taxiways, resulted in significant damage to the tail section of the smaller aircraft. Although no injuries were reported among the 277 passengers and crew, preliminary findings indicate that the pilot of the larger aircraft was momentarily distracted—likely due to efforts to monitor opposing traffic—thereby contributing to the mishap. Similar to case study I, this accident was caused by miscommunication between the tower controller and the pilot [14].
@@ 4.2.1 Link Travel Speed Parameters
For this case study, we look for a more realistic simulation with link speed distribution parameters obtained from ASDE-X from the Sherlock Data Warehouse (SDW). SDW is a comprehensive big data system developed specifically to support air traffic management (ATM) research, which collects raw data from multiple trusted sources such as the FAA and the National Oceanic Atmospheric Administration (NOAA). Data sources include flight plans, flight tracks from Air Route Traffic Control Centers (ARTCCs) and TRACON facilities, as well as meteorological such as wind, temperature, pressure, and precipitation from NOAA’s Rapid Refresh (RR) system, along with convective weather details (e.g., echo tops) from the Convective Integrated Weather Service (CIWS) and FAA SWIM. The raw input data are parsed and process to ensure that the data are reliable, queryable, and ready for further analysis. SDW is a central resource for researchers who need to analyze complex datasets and derive insights for improving air traffic operations and safety [118, 119, 120, 121].
Figure 9 provides the node-link graph layout of KATL, where several arrival flight trajectories from SDW ASDE-X are layered on top of it. According to [15], taxiways are further divided to crossover taxiways, apron taxiways, bypass taxiways, etc. For the sake of simplicity, the detailed classification of different taxiway types is not given here. Similar to previous work [100], we adopt the open source high performance geospatial data processing tool, Apache Sedona, to estimate the aircraft taxi speed at each node link.
Figure 10 and Figure 11 are the data analysis study results from the nodes of interest (i.e., overlapping nodes for both aircraft) from several days of ASDE-X. From Figure 10(a) and Figure 10(b), we show that Log-normal assumptions on link speed distributions are valid. Figure 10(c) and Figure 10(d) are the Quantile-Quantile (QQ) plot to compare the distribution of the data samples with assumed distributions. Moreover, the Kolmogorov-Smirnov (KS) test statistics are given, and the null hypothesis is that the sample follows a desired distribution, where a p-value greater or equal to are considered significant since the there is no significant evidence against it. The log-normal assumption is valid and convenient because it ensures positive speeds and provides tractable transformations for travel time. However, Figure 11(a) and Figure 11(b) show that at some links, the link speed distributions follow both normal and log-normal assumptions (i.e., p-value greater than 0.05), while normal distributions are better fits. To properly adapt to the risk formulation in Section 3.2.1, the normal distribution assumption of link speed are better used as truncated normal distributions and modify the spatial integral accordingly, or simply use log-normal across all links.
@@
Furthermore, a study on the impact of weight class to taxi speed is briefly conducted, to understand the impact of aircraft weight class to taxi speed. The weight class is derived into based on the FAA Order JO 7360.1E [122]. As shown in Table 3, Analysis of Variance (ANOVA) [123] and the Kruskal-Wallis [124] test are considered to study the impact of aircraft weight class both parametrically and non-parametrically. ANOVA is a statistical test used to determine whether there are significant differences between the means of two or more independent groups. It assumes that the data follows a normal distribution and that variances are equal across groups. The Kruskal-Wallis test is a non-parametric alternative to ANOVA, used when data does not meet normality or equal variance assumptions. Instead of comparing means, it ranks the data and compares the distributions across groups. The results for Link indicate that weight class influences taxi speed, while the test results from link is saying no strong evidence that weight class affects taxi speed. This suggests that the effect of weight class on taxi speed might be location-dependent or influenced by other factors like taxiway geometry, congestion, or operational procedures. To confirm a general relationship, further analysis (e.g., post-hoc tests for specific weight class differences, additional links) is needed and is listed as a major future study.
4.2.2 NER Results
@@
Once the real-world travel time parameters are obtained. We conduct the simulation and risk calculation as in Section 4.1. Similarly, Table 4 shows the NER model output as the guidance for taxiplan generation of the second simulation, where Figure 8 lists the progression of the taxiway collision case study. The associate risk score along each nodes and links are visualized in Figure 13. It is worth pointing out that only the nodes that are the overlaps of the two generated taxiplans are considered as potential collision spots, with a risk score.
4.3 Risk Visualization
These case studies highlight the inherent trade-offs between maintaining high operational throughput and ensuring safety in congested airport environments. They underscore the critical need for robust communication protocols, enhanced situational awareness, and the integration of advanced safety technologies. Our framework, which combines real-time data analysis with error detection mechanisms, aims to address these challenges and contribute to the reduction of runway and taxiway incidents in the future.

### v2

@@ 3.2 Airport Surface Collision Risk Model
Mathematically, the collision risk is calculated based on the joint probability that each aircraft reaches the same area at the same timestamp of the node-link airport layout graph. The location of a potential collision can be the intersection of runways, taxiways, taxilanes, or the combination of any two. The collision risk model is defined as the joint distribution of the time overlap for two aircraft occupying the same area around a node in the airport layout map. In this section, we provide the full probabilistic formulation of collision risk at a certain node in the node-link graph.
@@ 3.2.1 Total Travel Time Modeling
The -th aircraft travels a total of taxiway links until reaching the area of interest (i.e., potential collision spot), where the total travel time is given by . We assume each taxiway link has an associated distance and a taxi speed that is log-normally distributed with parameters and , which is
@@
In practice, we approximate for any time by either Monte Carlo Simulations or Moment-Matching Approximations. For the convolution of log-normal distributions with moderate variance and , the Fenton-Wilkinson approach provides a feasible solution to directly match the first two moments, and is widely-adopted as the approximated analytical solution of log-normal sums in various fields [117, 118, 119]. Fenton found that the sum of several independent log-normal variables can be reasonably approximated by another log-normal. Under this approximation, the mean of the route travel time equals the sum of link means, and the variance equals the sum of all link variances and covariance terms. These matched moments define the parameters of an approximate log-normal for the route [117]. Furthermore, [89] note that this log-normal approximation is computationally efficient compared to brute-force convolution or simulation, with only a modest loss of accuracy. That is, we are looking for the parameters of an approximate distribution of , where . That is,
and the associated cumulative density function (CDF) is as,
@@
3.2.2 Spatiotemporal Risk Formulation
As reviewed in Section 2.2, the collision between two moving aircraft on the surface is quantified by the probability of conflicts between aircraft by evaluating random deviations in position and speed. This also guides our spatiotemporal risk formulation. In our formulation, we assume that a collision occurs when two interchangeable aircraft satisfy the following two conditions,
@@ 3.2.2 Spatiotemporal Risk Formulation
We derive an expression to approximate the probability that a collision occurs at at any time . We start with a joint probability distribution for the spatiotemporal state of the aircraft. Define be the conditional PDF of aircraft 1’s arrival time at location , be the PDF of aircraft 2’s arrival time at location , and be the PDF describing aircraft 2’s spatial position at time .
@@
Approximating the velocity at collision as independent of and , a change of variables from space to time near can be made using the expected inverse of the speed at the point of collision,
@@
We prove Equation 16 using the joint density of the position, arrival time, and velocity and the relation . We assume that the aircraft velocity distribution is independent and that the time distribution is independent and constant in time such that . The position distribution is dependent on the time and velocity.
We first use Bayes’ rule for conditional distributions to relate the conditional position-velocity distribution to the conditional arrival time distribution.
We compute the distribution using the integral of joint probability over time and velocity. The relation is utilized for a change of variables from position to time. We also utilize the fact that is constant. The remaining integral results in the term.
The position distribution is plugged into the (18) and the constant time distribution terms cancel out.
@@
Additionally, the approximation error analysis of the proposed formulation is given in Appendix A.
3.2.3 Petri-Net Formulation
As an alternative to the proposed spatiotemporal risk formulation, we also provide a Petri-Net formulation for aircraft surface risk calculation. We follow the flow of [13] and propose the following formulation.
Following the previously defined notation, let be the aircraft index. Let denote the random arrival time of aircraft at the node (i.e., collision spot). Similarly, a graph-theoretic node is a point set of measure zero. Hence instantaneous occupancy at is zero in continuous time unless we introduce a small temporal window. We therefore define an operational coincidence window around the instant of arrival and the corresponding node-time occupancy indicator as,
Then the Petri-net node co-occupancy probability (within small ) is,
Lemma 3.
We realize Equation 28 here. Let denote the random arrival time of aircraft at node and define . Write and .
| (conditioning on ) | ||||||
| (definition of ) | ||||||
| (mean value theorem) | ||||||
where .
where is the value at zero of the PDF of in the FW formulation. is the probability the two arrivals fall within a coincidence window at the node. Further recall the FW-based risk probability at the same node with spatial capture radius is,
This is derived by converting the temporal overlap into a spatial hit via the (small) window and the expected inverse speed near . In the link model, using the incoming link’s .
Comparing the two expressions shows they share the same term . Eliminating gives the proportionality,
If we choose to match a temporal resolution (e.g., surveillance update or controller time bin) and pick as the spatial resolution (e.g., the half wingspan of the aircraft with a safety buffer), then the factor converts the PN node coincidence probability into the FW risk probability.
In summary, Petri-Net and FW formulations are tightly linked through the common overlap term where PN gives a temporal probability at the node. FW multiplies by a space–time factor to yield a collision risk probability at the same node. They are both probabilities. is a probability of temporal coincidence within at the node, which answers the question of Do the two arrivals occur within seconds of each other at the node? is a probability of a collision event at the node given a spatial capture radius .
3.2.4 Real-Time Risk Assessment
We enable real-time risk assessment capability of the node-based framework by inserting, at uniform temporal intervals (i.e., s), a virtual node at the instantaneous position of each aircraft along its current link. Let aircraft be on link at time , with link length and lognormal speed as in Section 3.2.1. Denote the residual distance on the current link as the remaining arc-length from the aircraft’s instantaneous position to the next topological node as,
Per the transformation in Equation 4, the residual time on the current link is,
with and obtained from the same moment formulas cited below Equation 4 by replacing with .
For any overlapping node downstream of the current position, define the residual route at time as the concatenation of the residual portion at and all full links from the next node up to . The residual travel time to is then a sum of link-times of the lognormal type in Equation 4. Its mean and variance are computed by summing the corresponding link moments (using the residual moments for the partial link and the original moments for full links). The route-level lognormal approximation,
is again obtained by the Fenton–Wilkinson moment-matching as in Equation 10. Moreover, the calculations of and remain the same as in Section 3.2.2.
The above construction is repeated for every and for every overlapping node . The per-node risk time series remain negligible until the aircraft converge near the terminal conflict node, at which point the overlap density and the associated risk rise sharply.
@@ 4 Case Studies
In this section, we present three illustrative case studies to demonstrate our proposed framework. The first case study examines the 2024 Haneda Airport runway collision in Section 4.1, the second focuses on the 2024 KATL taxiway collision accident in Section 4.2, and a reconstruction of the 1977 Tenerife airport disaster in Section 4.3. These accidents underscore the complex interplay of human error, communication misunderstandings, and operational failures. However, we use the assumed values for the link travel time speed distribution parameters of the first case study, based on recommended values mentioned in the FAA airport design recommendations, 150/5300-13B, as in [15]. For the second study, we obtain the ground movement data (i.e., ASDE-X) from Sherlock and use the link travel time speed distribution parameters from the real-world data. Details are given in Section 4.2.1. For the last one, we adopt the inferred taxi speed parameters based on the timestamp given from the literature [120].
4.1 Case I: Haneda Airport Runway Incursion
The Haneda Airport collision occurred on January 2, 2024, when Japan Airlines Flight 516 (Airbus A350) collided with a Japan Coast Guard DHC-8-315Q aircraft on the runway. The Coast Guard aircraft was stationed on the runway to deliver relief supplies following the Noto Peninsula earthquake. Despite the dramatic collision and ensuing fire, all 379 occupants aboard the A350 were successfully evacuated, while five of the six crew members on the smaller aircraft lost their lives. Investigations have attributed the accident primarily to miscommunication and human error where the Coast Guard pilot misinterpreted air traffic control instructions and mistakenly believed he had clearance to enter the runway. The communication transcripts ahead of the occurrence of the accident were released and tested against our developed learning model as in Figure 7(a) and Figure 7(b). It is clear that JA722A was given clearance to hold at C5 and should wait for departure clearance. However, the JA722A goes to the runway and misunderstood as the first to take off at runway 34R. Our ATC Rule-Enhanced Learning model processed the timeline of clearance given from the controller and the state of each aircraft and listed in Table 3.
@@ 4.1 Case I: Haneda Airport Runway Incursion
Based on Table 3, we build the surface movement simulation environment with the airport node-link graph layout obtained from NASA FACET [23]. We simulate the occurrence of the accident from the timestamp where the last clearance was given to JA722A from the ATCo in Figure 9. The collision happened at the timestamp of the lower left figure. The simulation takes three inputs for each link, the link distance , the link travel speed parameters . was calculated based on the location of the node on the node-link graph, with the speed parameters assumed to be if the link is between runway nodes, and between runway and taxiways. For links between two taxiway nodes, we assume the speed follows , and for all other scenarios (taxilanes, ramps, etc). These speed assumptions are based on the recommended values mentioned in the FAA airport design recommendations, 150/5300-13B, as in [15]. A detailed explanation of taxiways, runways, and terminology is given in Figure 11.
Table 3 shows that the NER model is able to capture the destination runway for both Japan Air 516 and Delta 276, where the key information such as callsign, aircraft state and also retrieved. The destination node from the node-link graph are correlated by either, (a) the entry node of the given runway; (b) the similarity between the destination node name and the entities from pilot-ATC communication transcripts. Table 3 visualizes several timestamps of the simulation, based on the assumed speed distribution parameters. Incorporating the information retrieved from the NER model, this simulation successfully replicates the occurrence of the Haneda runway incursion accident.
We provide the real-time risk at each overlapping nodes across the entire simulation time in Figure 10. In this scenario, these are nodes Rwy_03_006 to Rwy_03_011. The collision happens at the first overlapping node while the collision at other nodes remain subtle. For risk probability tolerance threshold at 0.05, the FW formulation provides a warning at around 41 seconds while the PN formulation fails to show the warning as in Figure 10(a). For lower risk tolerance at 0.01, both methods show warning but FW shows larger lead time compared with the PN approach as in Figure 10(b).
4.2 Case II: KATL Taxiway Collision
In the second case study, we investigate the taxiway collision occurred at KATL on September 10, 2024, where two Delta Air Lines aircraft collided while taxiing. An Airbus A350, preparing for an international departure to Tokyo, struck the tail of an Endeavor Air Bombardier CRJ-900, which was scheduled for a domestic flight to Lafayette, Louisiana. The collision, which occurred at the intersection of two taxiways, resulted in significant damage to the tail section of the smaller aircraft. Although no injuries were reported among the 277 passengers and crew, preliminary findings indicate that the pilot of the larger aircraft was momentarily distracted—likely due to efforts to monitor opposing traffic, thereby contributing to the mishap. Similar to case study I, this accident was caused by miscommunication between the tower controller and the pilot [14].
@@ 4.2.1 Link Travel Speed Parameters
For this case study, we look for a more realistic simulation with link speed distribution parameters obtained from ASDE-X from the Sherlock Data Warehouse (SDW). SDW is a comprehensive big data system developed specifically to support air traffic management (ATM) research, which collects raw data from multiple trusted sources such as the FAA and the National Oceanic and Atmospheric Administration (NOAA). Data sources include flight plans, flight tracks from Air Route Traffic Control Centers (ARTCCs) and TRACON facilities, as well as meteorological such as wind, temperature, pressure, and precipitation from NOAA’s Rapid Refresh (RR) system, along with convective weather details (e.g., echo tops) from the Convective Integrated Weather Service (CIWS) and FAA SWIM. The raw input data are parsed and processed to ensure that the data are reliable, queryable, and ready for further analysis. SDW is a central resource for researchers who need to analyze complex datasets and derive insights for improving air traffic operations and safety [122, 123, 124, 125].
Figure 11 provides the node-link graph layout of KATL, where several arrival flight trajectories from SDW ASDE-X are layered on top of it. According to [15], taxiways are further divided to crossover taxiways, apron taxiways, bypass taxiways, etc. For the sake of simplicity, the detailed classification of different taxiway types is not given here. Similar to previous work [102], we adopt the open source high performance geospatial data processing tool, Apache Sedona, to calculate the aircraft taxi speed at each node link.
Figure 12 and Figure 13 are the data analysis study results from the nodes of interest (i.e., overlapping nodes for both aircraft) from several days of ASDE-X. From Figure 12(a) and Figure 12(b), we show that Log-normal assumptions on link speed distributions are valid. Figure 12(c) and Figure 12(d) are the Quantile-Quantile (QQ) plot to compare the distribution of the data samples with assumed distributions. Moreover, the Kolmogorov-Smirnov (KS) test statistics are given, and the null hypothesis is that the sample follows a desired distribution, where a p-value greater or equal to are considered significant since there is no significant evidence against it. The log-normal assumption is valid and convenient because it ensures positive speeds and provides tractable transformations for travel time. However, Figure 13(a) and Figure 13(b) show that at some links, the link speed distributions follow both normal and log-normal assumptions (i.e., p-value greater than 0.05), while normal distributions are better fits. To properly adapt to the risk formulation in Section 3.2.1, the normal distribution assumptions of link speed are better used as truncated normal distributions and modify the spatial integral accordingly, or simply use log-normal across all links.
@@
Furthermore, a study on the impact of weight class on taxi speed is briefly conducted, to understand the impact of aircraft weight class to taxi speed. The weight class is derived based on the FAA Order JO 7360.1E [126]. As shown in Table 4, Analysis of Variance (ANOVA) [127] and the Kruskal-Wallis [128] test are considered to study the impact of aircraft weight class both parametrically and non-parametrically. ANOVA is a statistical test used to determine whether there are significant differences between the means of two or more independent groups. It assumes that the data follows a normal distribution and that variances are equal across groups. The Kruskal-Wallis test is a non-parametric alternative to ANOVA, used when data does not meet normality or equal variance assumptions. Instead of comparing means, it ranks the data and compares the distributions across groups. The results for Link indicate that weight class influences taxi speed, while the test results from link is saying no strong evidence that weight class affects taxi speed. This suggests that the effect of weight class on taxi speed might be location-dependent or influenced by other factors like taxiway geometry, congestion, or operational procedures. To confirm a general relationship, further analysis (e.g., post-hoc tests for specific weight class differences, additional links) is needed and is listed as a major future study.
@@
Once the real-world travel time parameters are obtained. We conduct the simulation and risk calculation as in Section 4.1. Similarly, Table 5 shows the NER model output as the guidance for taxiplan generation of the second simulation, where Figure 9 lists the progression of the taxiway collision case study. The associated risk score along each node and links are visualized in Figure 15. It is worth pointing out that only the nodes that are the overlaps of the two generated taxiplans are considered as potential collision spots, with a risk score.
Again, we report the real-time risk curves for the overlapping nodes Txy_E_004, Txy_E_003, and Txy_E_002 in Figure 15. The simulation runs for around 120 seconds. Similar to case 1, only the terminal conflict node (Txy_E_002) exhibits significant risk, while upstream nodes remain negligible. For the 0.05 threshold, FW crosses at 113 s while PN crosses at 117 s. At the lower threshold of 0.01, FW triggers a warning earlier at 110 s, and PN follows at 113 s.
This result highlights how risk accumulates only when both aircraft are on converging trajectories toward the same conflict node. The model’s selectivity ensures that nodes without simultaneous presence remain at zero risk, avoiding false positives. The FW lead time advantage is once again evident, providing a 3–4 second earlier warning compared to PN. Although the absolute time window is shorter than in the previous case, the relative lead time difference remains operationally significant, offering a larger buffer for mitigation actions.
4.3 Case III: Tenerife Runway Collision
The Los Rodeos airport disaster at Tenerife happened on March 27, 1977, and it is one of the deadliest accidents in aviation history, where KLM Flight 4805 and Pan Am Flight 1736 collided on the runway and killed 583 people [120, 129, 130]. Figure 16 provides a debriefing of the accident, while [131] provides the open-source transcribed communication recordings.
The accident stemmed from a series of external disruptions, human errors, and communication failures that converged under poor weather conditions, but the chain of miscommunication proved decisive. While KLM awaited clearance, the KLM pilot advanced throttles before explicit takeoff permission was granted [132]. KLM First Officer’s radio call, we are now at takeoff, overlapped with Pan Am’s transmission that they were still on the runway, producing a heterodyne effect that blocked critical ATC instructions [130]. KLM only heard OK from the tower, misinterpreting it as clearance. Meanwhile, Pan Am’s were unsure about where taxiway exit C3 is as instructed by the tower, and they continued to C4. Although Pan Am called that they were still on the runway, it was never received by KLM due to radio interference. This disaster reshaped civil aviation operations, as English proficiency requirements for ATC were reinforced and standard phraseology was adopted so that takeoff is used only when clearance is explicitly granted, with departure substituted in all other contexts [130].
| TIME | CALLSIGN | AC_STATE | DEST_RUNWAY | DESTINATION |
| 1658:25.7 | KLM 4805 | backtrack,takeoff | 30 | Rwy_12_006 |
| 1658:30.4 | KLM 4805 | taxi | 30 | Rwy_12_006 |
| 1658:47.4 | KLM 4805 | entering | 30 | Rwy_12_006 |
| 1658:55.3 | KLM 4805 | taxi | 30 | runway |
| 1659:28.4 | KLM 4805 | approach | 30 | Charlie 1(Rwy_12_001)? |
| 1701:57.0 | Clipper 1736 | |||
| 1702:03.6 | Clipper 1736 | taxi | 30 | runway? |
| 1702:16.4 | third (Rwy_12_003) | |||
| 1702:55.6 | KLM 4805 | pass | 30 | Charlie 4(Rwy_12_004) |
| 1705:44.6 | KLM 4805 | ready for takeoff | 30 | |
| 1706:09.6 | KLM 4805 | cleared,right turn | 30 | |
| 1706:12.25 | KLM 4805 | go | 30 | |
| 1706:20.08 | Clipper 1736 | taxi | 30 | runway |
| 1706:50.00 | KLM 4805 | collision | ||
| 1706:50.00 | Clipper 1736 | collision |
Table 6 shows the NER screening outputs. Due to the frequent use of non-standard ATC phraseology and frequent readbacks, the quality of the meta table is worse than the previous two case studies. The transcript reflects non-standard phraseology and period-specific practices (not uniformly adopted then), such as overlapping transmissions and ambiguous acknowledgments, which blur speaker identity, callsign boundaries, and clearance semantics. Thus, we perform several post-processing steps to further refine the output. In this table, assumed entities are filled and shown in bold. For instance, rows with a question mark indicate the entities are extracted from a question from the pilot to the controller.
Similar to the previous two case studies, we construct a simulation of the Tenerife accident using the node-link graph representation shown in Figure 17. For this case, we develop a custom node-link graph of Los Rodeos Airport (now Tenerife North Airport), since it is not available in NASA FACET. The graph is designed to replicate the 1977 airport layout and naming conventions to ensure consistency with historical conditions (i.e., the taxiway name changes from Charlie to Echo now).
Figure 18 visualizes the risk assessment results with the formulation in Section 3.2. Specifically, the dynamic risk is calculated with the approach introduced in Section 3.2.4, where the risk is repeatedly calculated at a fixed time interval. The real-time risk is shown at each overlapping node in Figure 18. In this scenario, overlapping nodes span Rwy_12_001 to Rwy_12_005, but the collision occurs at the final overlapping node Rwy_12_005. Throughout most of the 457 second simulation, the risk curves remain near zero. A sharp rise appears only in the final 10 seconds, when the two aircraft converge on the same segment. With a risk threshold of 0.05, FW triggers a warning at 450 s while PN triggers at 456 s. Under the more conservative threshold of 0.01, the warning occurs even earlier (448 s for FW, 452 s for PN).
These results emphasize two features. First, the model correctly localizes the risk to the true collision node, showing negligible risk at upstream intersections. Second, the FW formulation consistently provides a larger lead time compared to PN, in line with its scaling by aircraft geometry and speed. The steep terminal increase mirrors the operational reality: risk was negligible until both aircraft were committed to the final segment, at which point it rose rapidly to critical levels.